Project Status: 31 Jan 2007
Weather Status: Large amplitude blocking ridge still holds sway to the west of the HMT area. Dry conditions will continue for at least the next 5 days. Long range forecasts are indicating a general breakdown and shift east of this ridge some time in the 5-6 Feb time frame. Two models (Canadian and NOGAPS) have an event progged for 5 Feb. Other models GFS, ECMWF, and most GFS ensemble members are delaying this event. NWS is keeping precipitation out of the extended range forecasts at this time. Details of the regime change are still very much unknown. Much of the available forecasts and other tools are highly divergent at this time. So...no IOP expected in the next 5 days ( to 5 Feb), but potential exists for an event between 6-10 Feb.
System Status: Report from WR and SAC indicates that owing to personnel problems the OAK supplemental soundings will be unavailable for at least 4 weeks. Best scenario would see these available again in March. HMT and WR staff are exploring options to carry these un-used sondes over to next year's operations.